Calculate q, w, delta E and delta H for the process in which 2.0 moles of N2O gas are cooled from 165 degrees Celsius to 55 degrees Celsius at a constant pressure of 5 atm.
1
Expert's answer
2012-02-23T09:36:42-0500
delta E=0
delta H = Cp delta T
delta U = Cv delta T + p delta V
q = delta H
w = p delta V

delta U = dq + dw
w = -p deltaV = -R delta T = 8.31*(165 - 55) = 872.55 kJ